kindly help me please i have to submit this question with in 6 hours i have tried a lot but wasn't able to.........
if N-1=2/3,THEN N+1=? OPTIONS ARE
(A)4/3
(B)5/3
(C)7/3
(D)8/3
(E)14/3

Answer by TimothyLamb(4379)

If:
N - 1 = 2/3
Then:
N + 1 = ?
---
N - 1 = 2/3
N = 2/3 + 1
N = 2/3 + 3/3
N = 5/3
---
N + 1 = ?
N + 1 = 5/3 + 1
N + 1 = 5/3 + 3/3
---
answer:
N + 1 = 8/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php